how many liters of a 25% acid solution must be added to 20 liters of a 45% acid solution to get a 35% acid solution?
Answer by checkley77(12844)

.45*20+.25X=.35(20+X)
9+.25X=7+.35X
.25X-.35X=7-9
-.1X=2
X=2/.1
X=20 LITERS OF 25% ACID IS USED.
PROOF:
.45*20+.25*20=.35(20+20)
9+5=.35*40
14=14